A company issues $25,000,000, 7.8%, 20-year bonds to yield 8% on January 1,2017. Interest is paid on June 30 and December 31. The proceeds from the bonds are $24,505,180. What is interest expense for 2018, using straight-line amortization?
Answer #1

Par value of bonds = $25,000,000

Cash proceeds from issue of bonds = $24,505,180

Discount on bonds payable = Par value of bonds - Cash proceeds from issue of bonds

= 25,000,000-24,505,180

= $494,820

Semi annual interest payment = Par value of bonds x Interest rate x 6/12

= 25,000,000 x 7.8% x 6/12

= $975,000

Semi annual amortization of bonds discount = Discount on bonds payable /40

= 494,820/40

= $12,370.50

Semi annual interest expense = Semi annual interest payment + Semi annual amortization of bonds discount

= 975,000+12,370.50

= $987,370.50

Annual interest expense = Semi annual interest expense x 2

= 987,370.50 x 2

= $1,974,741

Interest expense for 2018 = $1,974.741

Kindly give a positive rating if you are satisfied with this solution and please ask if you have any query.

Thanks